OpenStudy (ebydaqueen19):

HELP HELP HELP STUCK ON THIS ONE 5b-{7[2(3b-2)-(4b+9)]-2}

OpenStudy (ebydaqueen19):

this is what i did

OpenStudy (ebydaqueen19):

lemme show u

OpenStudy (andijo76):

are you factoring or simplify

OpenStudy (ebydaqueen19):

oh never mind

OpenStudy (ebydaqueen19):

that was the wrong problem

OpenStudy (ebydaqueen19):

i dint get this one..wait a minute lemme type it

OpenStudy (ebydaqueen19):

this one u solve it regularly

OpenStudy (ebydaqueen19):

p-(p+4)=4(p-1)+2p

OpenStudy (ebydaqueen19):

this is how i did it

OpenStudy (ebydaqueen19):

p-p+4=4p-4+2p 4=6p-4 -6p+4=-4 -6p=-8 p=4/3

OpenStudy (ebydaqueen19):

correct answer is 0 and i got 4/3 :/

OpenStudy (ebydaqueen19):

could u let me know what step i did wrong

OpenStudy (andijo76):

yeah the correct answer is 0 that is what i see

OpenStudy (andijo76):

i found a website with the step by step if you want i can paste it for you on here ( the answer it gave me)

OpenStudy (ebydaqueen19):

yes please :)

OpenStudy (andijo76):

p−(p+4)=4(p−1)+2p Multiply 4 by each term inside the parentheses. More Detail Multiply 4 by each term inside the parentheses (p−1). p−(p+4)=4(p)+4(−1)+2p Multiply 4 by the p inside the parentheses. p−(p+4)=4⋅p+4(−1)+2p Multiply 4 by p to get 4p. p−(p+4)=4p+4(−1)+2p Multiply 4 by the −1 inside the parentheses. p−(p+4)=4p−4⋅1+2p Multiply −4 by 1 to get −4. p−(p+4)=4p−4+2p p−(p+4)=4p−4+2p Since 4p and 2p are like terms, add 2p to 4p to get 6p. p−(p+4)=6p−4 Multiply −1 by each term inside the parentheses. Multiply −1 by each term inside the parentheses (p+4). p−(p)−(4)=6p−4 Multiply −1 by the p inside the parentheses. p−p−(4)=6p−4 Multiply −1 by the 4 inside the parentheses. p−p−4=6p−4 p−p−4=6p−4 Since p and −p are like terms, add −p to p to get 0. 0−4=6p−4 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. −4=6p−4 Since p is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 6p−4=−4 Since −4 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 4 to both sides. 6p=4−4 Subtract 4 from 4 to get 0. 6p=0 Divide each term in the equation by 6. 6p6=06 Simplify the right-hand side of the equation. Cancel the common factor of 6 in 6p6. 1/ 6 p 6 =06 Reduce the expression 6p6 by removing a factor of 6 from the numerator and denominator. p=0/6 p=0/6 Simplify the equation. p=0

OpenStudy (andijo76):

i hope you understand this sometimes it does not paste right

OpenStudy (ebydaqueen19):

thx

OpenStudy (ebydaqueen19):

what is the website name?